Evaporation of water from our skin is a very important mechanism for controlling body temperature.

In a 4.5-h intense cycling race, a cyclist can loses 9.5 kg of water through perspiration. Estimate how much heat the cyclist loses in the process.
Answer #1

considering normal body temperature to be 37 deg C .

first energy needed to raise temperature to 100 deg C.

Q1 = m c deltT = 9.5 x 4186 x (100 - 37) = 2505321 J

energy needed to evaporation ,

Q2 = mL = 9.5 x 2256 x 10^3 = 21432000 J


total heat lost = Q1 + Q2 = 23927321 J